OpenStudy (anonymous):

You are going to fence an area that is in the shape of a right triangle. The two legs measure 7.8 m and 5.78 m. Estimate the meters of fencing you will need for the longest side, using whole meters and square root. 35 m 48 m 10 m 24 m

OpenStudy (anonymous):

24 m ( 5.78 taken as 6, 7.8 taken as 8, hence biggest side sqrt(8^2 + 6 ^ 2) = 10) Hence perimeter in total = 10 + 8 + 6 = 24